in 2 years andy will be as old as joe is now. in 2 years joe will be 1 1/2 times as old as andy. find their ages now
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Let joe's age be x.
Andy's age will be x-2 now. ( because after two years andy will be joe's age.)
After two years the relation of their ages will be
x+2 = 3/2 * x ( joe is one and half times andy's age.)
2(x+2)=3
2x+4=3x
3x-2x=4
x=4 Joe's age, 4 years.
2years- Andy's age

Answer by richwmiller(17219)

j=joe now
a=andy now
j+2=1.5(a+2)
a+2=j
a+2+2=1.5(a+2)
a+4=1.5a+1.2*2
a+4=1.5a+3
1=.5a
2=a
4=j